# Unlocking Text Capabilities in Vision Models

### v2

@@ Abstract
Visual classifiers provide high-dimensional feature representations that are challenging to interpret and analyze. Text, in contrast, provides a more expressive and human-friendly interpretable medium for understanding and analyzing model behavior. We propose a simple, yet powerful method for reformulating any pretrained visual classifier so that it can be queried with free-form text without compromising its original performance. Our approach is label-free, data and compute-efficient, and is trained to preserve the underlying classifier’s distribution and decision-making processes. Our method unlocks several zero-shot text interpretability applications for any visual classifier. We apply our method on 40 visual classifiers and demonstrate two primary applications: 1) building both label-free and zero-shot concept bottleneck models and therefore converting any visual classifier to be inherently-interpretable and 2) zero-shot decoding of visual features into natural language sentences. In both tasks we establish new state-of-the-art results, outperforming existing works and surpassing CLIP-based baselines with ImageNet-only trained classifiers, while using up to 400$\times$ fewer images and 400,000$\times$ less text during training.
@@ 1 Introduction
Visual classifiers provide dense, high-dimensional visual feature vectors that are difficult to interpret directly. Attribution maps—the standard way to interpret visual classifiers—only highlight the image region most influential to a prediction, offering a high-level spatial insight. They reveal where the key features lie, but not what those features actually mean or represent. By contrast, text is significantly more expressive, more precise and human-friendly, and it enables a new suite of text interpretability tasks  [1, 2, 3]. This raises a key question: How can we query a purely visual classifier with free-form text? Performing this allows us to have a text interface that can be used to query visual classifiers. Contrastive Language-Image Pre-Training (CLIP)-based vision–language models answer this by learning a shared embedding space [4], enabling any text to query the visual model directly. Yet in many real-world scenarios a high-performing, task-specific legacy model already exists and typically outperforms zero-shot vision-language models [5]. Retraining such a specialist on a large image–text corpus following the CLIP approach or the GPT-like approach is impractical, in terms of computational cost and need of a huge amount of image-text data. Moreover, this approach alters the decision-making process and distribution of the legacy model. As a result, most text-based interpretability techniques remain confined to CLIP-like models, leaving specialist vision models without a language-driven interface.
Figure 1: (a) Text interpretability is typically limited to models trained to share a joint embedding space such as CLIP-based ones. (b) Our approach removes this constraint, enabling any visual classifier to be queried through any arbitrary (unaligned) text encoder in a highly data- and compute-efficient, label-free manner, while preserving the classifier’s original distribution and reasoning process and thus offering text-based interpretability of the original classifier. The values here represent the cosine similarity between the text and the visual features $f$.
@@
In this work, we lift this limitation by proposing a method to reformulate any frozen visual classifier such that it can be accessed via free-form text (Figure 1). Our method is characterized by four important properties: First, it is efficient; it is inexpensive to train and can be performed on any standard hardware, regardless of the size of the original classifier. The number of data points is also significantly reduced compared to CLIP-based approaches. Secondly, it is label-free, no labels are required to achieve this formulation. Thirdly, it is explicitly trained to preserve the original distribution and reasoning process of the classifier. Finally, our method is applicable to any vision architecture, whether convolutional-based, transformer-based or hybrid.
A visual classifier assigns an image to a specific category from a predefined set of discrete class labels. In ImageNet-1K [6], this set contains $1,000$ class labels. Originally, these discrete class labels correspond to class names in text format. For example, in ImageNet-trained models, the discrete label $1$ corresponds to the class goldfish. These classes are typically discretized to facilitate training with cross-entropy. However, when the textual class names are embedded into vector representations (e.g, using a text encoder or word embedding model), they provide semantic information. Specifically, these embeddings reside in a continuous space where nearby vectors capture related concepts. For the “goldfish" example, neighboring vectors might include “freshwater", “fins" and “orange". Our method learns to map images into this text embedding space using only class names, thus associating both the class name and its surrounding semantic space with the image. This is accomplished through a trainable multilayer perceptron (MLP) that projects the visual features into the text embedding space, and is explicitly trained to output the same distribution scores across all classes, as the original classifier’s class distribution. This is done while while keeping both the visual and textual encoders frozen. By using solely the class names without any supplementary information, we can learn a semantically meaningful image-text space. This allows us to query the visual classifier with any text query beyond the class names.
We demonstrate the effectiveness of our method with two primary applications. 1) We build label-free concept bottleneck models for any pretrained classifier (not being restricted to CLIP) while obtaining the linear probe classifier in a zero-shot manner (we do not train a linear probe on the concept activations), also ensuring that the newly built linear classifier shares the same reasoning process with the original classifier. 2) We show that our method can decode visual features of any classifier into natural language in a zero-shot manner, without requiring training on image-text data such as image-caption pairs, while also showing the generalization of our method to datasets beyond ImageNet.
In summary, our contributions are as follows: (i) We propose an efficient, label-free method to reformulate any visual classifier such that it can be queried with open-set text queries, maintaining its performance within a negligible margin (average of 0.2 points drop in accuracy) (ii) We demonstrate the effectiveness of our method with 40 different architectures through two applications: label-free zero-shot Concept Bottleneck Models (CBMs) for any visual classifier, and zero-shot decoding of visual features into natural language. (iii) On both these applications, our method sets new state-of-the-art results outperforming existing works, including those using CLIP as the visual encoder, while being trained with 400$\times$ less images and 400,000$\times$ less text data.
@@ 2 Related Work
While CLIP is the predominant approach to interpret vision models through language [2], there exist works that try to decode visual features of models beyond CLIP. DeVIL [7] and LIMBER [8] trains an autoregressive text generator to map visual features into image captions, leveraging annotated image-caption pairs as ground-truth data. Similarly, Natural Language Explanations (NLEs) [9, 10, 11] use annotated textual explanations in place of conventional captions. Notably, all these works (1) rely on annotated datasets and (2) explicitly train the generated text to align with what annotators want the visual features to describe, thereby altering the classifier’s original reasoning process. ZS-A2T [12] converts attention maps into natural language in a zero-shot manner using LLMs, but is constrained to vision-language models trained to learn a shared vision-language embedding space (e.g., through a contrastive objective); therefore, it cannot be applied to standalone visual classifiers. In contrast, our method decodes visual features of any pretrained classifier in a zero-shot manner without requiring any annotated data, and is explicitly trained to keep the model’s predictive distribution intact.
Text-to-Concept (T2C) [13] and [14] are the closest related works to ours. These works train a linear layer to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder. Since the linear layer maps features to the CLIP space, these works are strongly biased towards interpreting the CLIP model rather than the original classifier. As the ground-truth data for training the linear mapper are the features extracted from the CLIP vision encoder, this method is approximate to using the CLIP vision encoder directly to encode the image. Moreover, these works rely on CLIP supervision during training. In contrast, our approach is label-free and our mapping function is explicitly trained to preserve the classifier’s original distribution.
@@ 3 Method
Figure 2: An overview of our method for reformulating visual classifiers. (a) The process of training the MLP mapping between vision and text space. A PyTorch-like pseudocode for (a) in included in Section B of the Appendix. (b) The process of inference with the adapted visual classifier. The text encoder acts as a linear classifier weight generator in our re-formulated classification process. Refer to caption indicates that the module is frozen, while Refer to caption indicates trainable.
@@
Consider an image $I$ and a visual classifier $F$ composed of a visual feature extractor $F_{v}$ and a linear classifier $W$. Note that $F$ can be of any architecture. $F_{v}$ embeds $I$ into an $n-$dimensional feature vector $f\in\mathbb{R}^{n}$. That is, $f=F_{v}(I)$. The linear classifier $W\in\mathbb{R}^{n\times K}$ takes $f$ as input and outputs a probability distribution $o$ for the image across $K$ classes. That is, $o=\text{softmax}(f.W)\in\mathbb{R}^{K}$. For ImageNet-1K, $K=1000$. Consider also any off-the-shelf text encoder $T$ which takes in an input text $l$ and embeds it into a $m-$dimensional vector representation $u\in\mathbb{R}^{m}$. That is, $u=T(l)$. Note that $u$ and $f$ are not in the same space and can have a different number of dimensions, so we cannot query $f$ with the text $l$.
@@
We propose to learn a simple lightweight MLP mapping function that projects the visual features $f$ into the text embedding space of $T$, resulting in a new vector $\tilde{f}$. That is, $\tilde{f}=\text{MLP}(f)$, where $\tilde{f}\in\mathbb{R}^{m}$. Note that the visual encoder $F_{v}$, the linear classifier $W$, and the text encoder $T$ are all frozen; only the MLP is trainable, making our method efficient. We then take the textual class names of the $K$ classes, and convert each into a text prompt $l^{p}$, represented as: “an image of a {class}" where {class} represents the class name in text format. This results in $K$ textual prompts, each of which is encoded with $T$: $u_{i}=T(l^{p}_{i})$, $\forall i=1,\dots,K$. Stacking all the encoded prompts, we get a matrix $U\in\mathbb{R}^{K\times m}$. Here, $U$ acts as weights of the classification layer for our approach. We then calculate the cosine similarity111in the rest of this paper, we will omit the unit norm in cosine similarity to reduce clutter, and represent it with the dot product between each $u_{i}$ and the visual features $f$: $s_{i}=\tilde{f}.u_{i}$. Equivalently, this can be performed as a single vector-matrix multiplication: $S=\tilde{f}.U^{T}$, where $S\in\mathbb{R}^{K}$ represents the cosine similarity scores between the visual features and every text prompt $l^{p}_{i}$ representing a class. In other words, $S$ represents the classification logits of our approach.
@@
The most straightforward way to training the MLP mapper is to leverage the ground-truth labels from the dataset, aligning $S$ with the ground-truth distribution. However, this approach violates two key properties: (1) it necessitates annotated data, and (2) re-training the classifier alters its original distribution $o$, thereby changing the reasoning process of the classifier (i.e., how it maps visual features to class probabilities and makes predictions) and sacrificing text interpretability of the original classifier. Notably, the original soft probability distribution $o$ is a function of the linear classifier $W$, so $W$ cannot be ignored. We instead propose to align $S$ to the original distribution $o$ through the cross-entropy loss. For a single sample, the loss is given by:
@@
This task can be viewed as a knowledge distillation problem, except that we do not distill the knowledge of a bigger teacher model to a smaller student model, but distill the distribution of the original model to a reformulated way of classification. Note that the loss is equivalent to the KL divergence loss between $o$ and the predicted distribution, since the additional entropy term $H(o)$ that appears in the KL divergence loss is a constant that does not depend on the MLP parameters. Eq. 1 shows that our approach is label-free and steers the MLP to be faithful to the original classifier $F$, since it is explicitly trained for that purpose. Note also that this loss function naturally encodes the classifier’s relationship across all classes.
Our training approach is illustrated in Figure 2(a), and we provide a PyTorch-like pseudocode of it in Section B of the Appendix. It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions, concepts or hierarchies (see Section I in the Appendix for more information).
After training, the projected visual features and the text encoder features lie in the same space, and we can therefore query the visual features with any text by finding the alignment score between the encoded text and the visual features. In the case of image classification, the text queries remain the class prompts, and encoding them with the text encoder $T$ is equivalent to generating the weights of a linear classifier for the newly formulated classification task, defined as argmax($\tilde{f}.U^{T}$). This is shown in Figure 2(b).
4 Applications
In this section, we show how our method can be used to build text-based interpretability applications. At application, all model components (including the MLP) are frozen.
@@ 4.1 Zero-Shot Concept Bottleneck Models
Concept Bottleneck Models (CBMs) [1] are a class of inherently interpretable models and have recently attracted significant attention. They consist of two steps: (1) concept discovery, followed by (2) concepts-to-predictions. In step (1), the dense output features of a visual encoder are first mapped to textual concepts (e.g., words or short descriptions of objects) each with a score that represents the concept activation to the image. In step (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class such that a prediction can be interpreted as a linear sum of interpretable concepts rather than a linear sum of opaque features. Recently, Label-Free CBMs (LF-CBMs) use CLIP to perform the concept discovery step without annotated image-concept data. This is achieved either by using CLIP as a model to provide ground-truth image-concept similarities to train a layer that performs concept discovery [15], or by querying the image features from a set of predefined concepts within the CLIP space [16, 17] and using the cosine similarity scores between the concepts and the image as concept activations. However, there are major limitations in this existing line of CBMs. Firstly, LF-CBMs are restricted to CLIP-based models, as the concept discovery step would otherwise not be possible since it necessitates an image-text similarity model. Secondly, all CBM approaches to date require training a linear probe on the concept activations, which 1) hinders an on-the-fly ready-to-deploy CBM at test-time directly, and 2) requires training a new $W^{con}$ classifier from scratch for every different set of concepts, hindering flexibility to the choice of concepts.
Our method solves all the aforementioned issues, formulating zero-shot CBMs for any pretrained classifier. We remind readers from Section 3 that $U\in\mathbb{R}^{K\times m}$ is the output of the text encoder $T$ for the class prompts, which in essence represent the classification weights of the newly formulated classifier. We assume that we are given a large set of predefined textual concepts, denoted as $\mathcal{Z}$, and with cardinality $|\mathcal{Z}|=Z$. Following other works [18], and without loss of generality, we use the $Z=20K$ most common words in English [19] as our concept set. These are general concepts that are sufficiently expressive and represent world knowledge and are not tailored towards any specific dataset. To ensure that the concepts are meaningful, we apply a rigorous filtering procedure to the concept set. Specifically, we remove any terms that exactly match the target class name, as well as any constituent words that form the class name (for example, eliminating “tiger” and “shark” when the class name is “tiger shark”). In addition, we exclude terms corresponding to the parent and subparent classes (e.g., “fish” and “animal” for the class “tiger shark”), other species within the same category, and any synonyms of the target class name. This systematic filtering guarantees that the resulting concept set is free of terms that are overly similar or directly derived from the target classes.
@@
Figure 3: (a) Our approach to concept discovery in CBMs. (b) Decoding visual features into natural language using any pretrained language decoder (e.g., GPT-2). We apply prefix tuning while keeping the language decoder frozen, generating text that maximizes the similarity with the visual features. Shown for iteration $j=1$
We now show how to build the classifier $W^{con}$ that takes in the concept activations and outputs a distribution $S_{cn}$ over classes. We build $W^{con}$ in a zero-shot manner. Here, zero-shot indicates that no training is required to map concept activations to classes. Recall that both $U$ and $C$ are outputs of the text encoder $T$, and they are already in the same space. Therefore, we can build the weights of the classifier $W^{con}$ with a text-to-text search between the concepts and the class name. Specifically, we calculate the cosine similarity between the concept embeddings $C$ and the linear classifier $U$ to obtain the new weights for $W^{con}$. That is, we perform $C\cdot U^{T}\in\mathbb{R}^{Z\times K}$. Therefore, the weights of $W^{con}$ represent how similar the class name is to all the concepts. In total, the output distribution $S_{cn}$ of the CBM is defined as:
@@
From Eq. 2 we observe that we scale the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$). Notably, if the gram matrix is identity ($C^{T}C=I$), we get back our original feature-based classifier given by $\tilde{f}.U^{T}$. Therefore, to convert any classifier to a CBM, all we have to do is plug in the gram matrix in-between, making it a convenient way to directly switch to an inherently interpretable model. Eq. 2 also shows that we do not change the linear classifier $U$, we only scale it by the gram matrix of concepts. This means our CBMs preserve the basis reasoning process of the original classifier. By this, we obtain zero-shot CBMs that discover concepts and builds $W^{con}$, both in a training-free manner and for any classifier. Note that our method is flexible in the concept set. The concept set is chosen at test time directly, and the CBM is built directly on-the-fly.
4.2 Zero-Shot Decoding of Visual Features into Text
In this application, we aim to decode the visual feature vector $f$ for an image $I$, given by $f=F_{v}(I)$, into a natural language sentence. This offers a text-based interpretation of what the visual features contain. We adapt the method introduced in [3] for our purpose. Specifically, we first project the visual feature vector $f$ using the MLP to obtain $\tilde{f}$. That is, $\tilde{f}=\text{MLP}(f)$. Since $\tilde{f}$ is now in the same space as the text encoder $T$, we can measure its association to any encoded text. We utilize an off-the-shelf pretrained language decoder model (e.g., GPT-2), denoted as $G$, to generate open-ended text. We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning [20] to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$. Specifically, we attach a set of learnable “virtual" tokens to $G$. Denote the generated output text of $G$ for one iteration as $h^{j}$, where $j$ represents the iteration number. $h^{j}$ is encoded with the text encoder $T$, to yield a vector $y^{j}$, That is, $y^{j}=T(h^{j})$. As $y^{j}$ and $\tilde{f}$ are now in the same embedding space, we maximize the cosine similarity between them in order to update the learnable tokens. We perform this process for several iterations. An overview of this process is shown in Figure 3(b). As this process is not the core contribution of our work, we leave more details to Section H of the appendix, which also includes a detailed illustration in Figure 2.

### v3

# CLIP-Free, Label-Free, Zero-Shot Concept Bottleneck Models

@@ Abstract
Concept Bottleneck Models (CBMs) map dense, high-dimensional feature representations into a set of human-interpretable concepts which are then combined linearly to make a prediction. However, modern CBMs rely on the CLIP model to establish a mapping from dense feature representations to textual concepts, and it remains unclear how to design CBMs for models other than CLIP. Methods that do not use CLIP instead require manual, labor intensive annotation to associate feature representations with concepts. Furthermore, all CBMs necessitate training a linear classifier to map the extracted concepts to class labels. In this work, we lift all three limitations simultaneously by proposing a method that converts any frozen visual classifier into a CBM without requiring image-concept labels (label-free), without relying on the CLIP model (CLIP-free), and by deriving the linear classifier in a zero-shot manner. Our method is formulated by aligning the original classifier’s distribution (over discrete class indices) with its corresponding vision-language counterpart distribution derived from textual class names, while preserving the classifier’s performance. The approach requires no ground-truth image–class annotations, and is highly data-efficient and preserves the classifiers reasoning process. Applied and tested on over 40 visual classifiers, our resulting CLIP-free, zero-shot CBM sets a new state of the art, surpassing even supervised CLIP-based CBMs. Finally, we also show that our method can be used for zero-shot image captioning, outperforming existing methods based on CLIP, and achieving state-of-art results.
@@ 1 Introduction
Visual classifiers predict a class as a linear combination of dense, high‑dimensional visual feature vectors that are difficult to interpret by humans. Concept Bottleneck Models (CBMs) (pmlr-v119-koh20a) address this challenge by mapping these feature vectors into a set of human-interpretable concepts, each associated with an activation score (referred to as a concept activation). Predictions are then made as a linear combination of these concept activations. Initial CBMs required image–concept annotations to train the bottleneck layer that maps dense features to concepts. Modern CBMs (oikarinen2023labelfree; Yang2022LanguageIA; Panousis2023SparseLC; Rao2024DiscoverthenNameTC) overcome this limitation by leveraging the CLIP model (Radford2021LearningTV) as an annotation-free alternative. Since CLIP models map image and text into a shared embedding space, these approaches can query image features against a set of predefined textual concepts within that space and use cosine similarity scores to find a matching annotation. These approaches are commonly referred to as label-free CBMs.
Nevertheless, in many real-world scenarios a high‑performing, task‑specific legacy model already exists and typically outperforms zero‑shot CLIP models (zhang2024why). A natural question that then arises is: how to develop a label-free CBM for such legacy specialist models? Retraining such a specialist on a large image–text corpus following the CLIP approach is impractical, in terms of computational cost and need of a huge amount of image-text data. Furthermore, obtaining ground-truth image-concept annotations is time consuming and labor intensive. Finally, retraining this legacy model further alters its original decision-making process and distribution, which is typically not desired. As a result, CBMs remain confined to CLIP‑like models, leaving specialist vision models aside with no clear way to adapt them into the CBM framework.
@@
In this work, we lift this limitation by first proposing a method dubbed as TextUnlock, which aligns a distribution of a frozen visual classifier to its corresponding vision-language counterpart, without relying on CLIP. TextUnlock has four important properties: First, it is efficient; it is inexpensive to train and can be performed on any standard hardware, regardless of the size of the original classifier. The number of data points is also significantly reduced compared to CLIP-based approaches. Secondly, it is label-free, no labels are required to achieve this formulation. Thirdly, TextUnlock is trained to preserve the original distribution and reasoning process of the classifier, and does not compromise the classifier’s original performance (average of 0.2 points drop in accuracy). Finally, our method is applicable to any vision architecture, whether convolutional-based, transformer-based or hybrid. After the original classifier’s distribution is aligned to its vision-language counterpart distribution with TextUnlock, we simply query the transformed classifier’s image features against a set of predefined text concepts to obtain concept activations for our CBM, and then derive the concept-to-class classifier directly from the transformed classifier’s text source, without requiring any additional training but rather operating in a zero-shot manner (i.e., we do not train a linear probe to associate the concept activations to class labels).
In summary, our contributions are as follows: (i) We propose a method to convert any frozen visual classifier into a CBM without relying on the CLIP bottleneck (CLIP-free), neither annotated image-concept data (label-free) while also deriving the concept-to-class classifier in a zero-shot manner. (ii) We demonstrate the effectiveness of our method with 40 different architectures, along with extensive ablation studies and intervention results, and further show how our method can also be used to perform zero-shot image captioning. (iii) Our method sets new state-of-the-art results, outperforming existing works including supervised CLIP-based CBMs, despite being trained only on ImageNet-1K.
@@ 2 Related Work
Concept Bottleneck Models (CBMs): Concept Bottleneck Models (CBMs) (pmlr-v119-koh20a) are a class of inherently interpretable models that map dense representations into human-understandable concepts, and then make predictions based on a linear combination of these interpretable units rather than opaque features. Recently, Label-Free CBMs (LF-CBMs) have been introduced, which leverage CLIP to perform the feature-to-concept mapping without requiring annotated image–concept data. This can be achieved either by using CLIP to provide ground-truth image–concept similarity scores that train the bottleneck layer (oikarinen2023labelfree), or by querying image features against a predefined set of concepts in the CLIP space (Yang2022LanguageIA; Panousis2023SparseLC) and using the resulting cosine similarities as concept activations. Several follow-up works (Rao2024DiscoverthenNameTC; Knab2024DCBMDV; Benou2025ShowAT; yuksekgonul2023posthoc) adopt this CLIP-based pipeline. However, these approaches remain limited to CLIP-based classifiers and cannot be applied to other visual models. Moreover, they require training a linear probe to map concept activations to class predictions. Our work addresses both limitations: we propose a fully CLIP-free CBM and demonstrate that our method can be readily used to derive the linear probe in a fully zero-shot manner.
Transforming Visual Features to Text: There exist works that try to decode visual features of models beyond CLIP. DeVIL (Dani2023DeViLDV) and LIMBER (merullo2023linearly) train an autoregressive text generator to map visual features into image captions, leveraging annotated image-caption pairs as ground-truth data. Notably, all these works (1) rely on annotated datasets and (2) explicitly train the generated text to align with what annotators want the visual features to describe, thereby altering the classifier’s original reasoning process. ZS-A2T (Salewski2023ZeroshotTO) converts attention maps into natural language in a zero-shot manner using Large Language Models (LLMs), but is constrained to vision-language models trained to learn a shared vision-language embedding space (e.g., through a contrastive objective). Text-to-Concept (T2C) (Moayeri2023TextToConceptB) also trains a linear layer to map image features of any classifier into the CLIP vision encoder space, such that they can be interpreted via text using the CLIP text encoder. Notably, all these methods rely on the CLIP approach and/or its supervision, and they alter the classifier’s distribution by entirely discarding its output class distribution. In contrast, our method is CLIP-free, can be applied to any pretrained classifier without requiring any annotated data, and explicitly preserves the classifier’s predictive distribution.
@@ 3 Method
We first elaborate on our proposed TextUnlock method in Section 3.1, which will then allow us to design zero-shot concept bottleneck models, which we discuss in Section 3.2.
3.1 TextUnlock
A visual classifier assigns an image to a specific category from a predefined set of discrete class labels. For example, in ImageNet-1K (Deng2009ImageNetAL), this set contains $1,000$ class labels. Originally, these discrete class labels correspond to class names in text format. For example, in ImageNet-trained models, the discrete label $1$ corresponds to the class goldfish. These classes are typically discretized to facilitate training with cross-entropy. However, when the textual class names are used, they provide an advantage. Specifically, when the textual class names are embedded into vector representations (e.g., using a word embedding model or text encoder), they provide semantic information. Specifically, These embeddings reside in a continuous space where nearby vectors capture related semantic associations. For the “goldfish” example, neighboring vectors might include terms such as “freshwater”, “fins” and “orange”. Such associations can be viewed as high-level conceptual attributes that characterize the class. Our method learns to map images into this text embedding space using only class names, thus linking both the class name and its surrounding semantic associations with the image. This process thus naturally supports unseen words that are not part of the class names (e.g., concepts in the CBM). We accomplish this through a trainable multilayer perceptron (MLP) that projects the visual features into the text embedding space, and is explicitly trained to match its distribution with the original classifier’s class distribution. This is done while keeping both the visual and textual encoders frozen. By using solely the class names without any supplementary information, we can learn a semantically meaningful image-text space. This allows us to query the visual classifier with text queries beyond the class names (e.g., concepts) and obtain concept activations for CBMs.
@@ 3.1 TextUnlock
Consider an image $I$ and a visual classifier $F$ composed of a visual feature extractor $F_{v}$ and a linear classifier $W$. Note that $F$ can be of any architecture. $F_{v}$ embeds $I$ into an $n$-dimensional feature vector $f\in\mathbb{R}^{n}$. That is, $f=F_{v}(I)$. The linear classifier $W\in\mathbb{R}^{n\times K}$ takes $f$ as input and outputs a probability distribution $o$ for the image across $K$ classes. That is, $o=\text{softmax}(f.W)\in\mathbb{R}^{K}$. For ImageNet-1K, $K=1000$. Consider also any off-the-shelf text embedding model $T$ which takes in an input text $l$ and embeds it into a $m$-dimensional vector representation $u\in\mathbb{R}^{m}$. That is, $u=T(l)$. Note that $u$ and $f$ are not in the same space and can have a different number of dimensions, so we cannot query $f$ with the text $l$.
Figure 1: Overview of our proposed TextUnlock. (a) The process of training the MLP mapping between vision and text space with pseudocode given in Appendix Section A. (b) The process of inference with the adapted visual classifier. The text encoder acts as weight generator for a linear classifier. Refer to caption indicates that the module is frozen, while Refer to caption indicates trainable.
@@
We propose to learn a lightweight MLP mapping function that projects the visual features $f$ into the text embedding space of $T$, resulting in a new vector $\tilde{f}$. That is, $\tilde{f}=\text{MLP}(f)$, where $\tilde{f}\in\mathbb{R}^{m}$. Note that the visual encoder $F_{v}$, the linear classifier $W$, and the text encoder $T$ are all frozen; only the MLP is trainable, making our method parameter-efficient. We then take the textual class names of the $K$ classes, and convert each into a text prompt $l^{p}$, represented as: “an image of a {class}" where {class} represents the class name in text format. This results in $K$ textual prompts, each of which is encoded with $T$: $u_{i}=T(l^{p}_{i})$, $\forall i=1,\dots,K$. Stacking all the encoded prompts, we get a matrix $U\in\mathbb{R}^{K\times m}$. Here, $U$ acts as weights of the classification layer for our approach. We then calculate the cosine similarity111in the rest of this paper, we will omit the unit norm in cosine similarity to reduce clutter, and represent it with the dot product. between each $u_{i}$ and the visual features $f$: $s_{i}=\tilde{f}.u_{i}$. Equivalently, this can be performed as a single vector-matrix multiplication: $S=\tilde{f}.U^{T}$, where $S\in\mathbb{R}^{K}$ represents the cosine similarity scores between the visual features and every text prompt $l^{p}_{i}$ representing a class. In other words, $S$ represents the classification logits of our approach.
@@
The most straightforward way to training the MLP is to leverage the ground-truth labels from the dataset, aligning $S$ with the ground-truth distribution. However, this approach violates two key desiderata: (1) it necessitates annotated data, and (2) re-training the legacy classifier alters its original decision distribution $o$, thereby changing the reasoning process of the classifier (i.e., how it maps visual features to class probabilities and makes predictions). Notably, the original soft probability distribution $o$ is a function of the linear classifier $W$, so $W$ cannot be ignored. We instead propose to align $S$ to the original decision distribution $o$ through cross-entropy loss. For a single sample, the loss is given by
@@
This task can be viewed as a knowledge distillation problem, except that we do not distill the knowledge of a bigger teacher model to a smaller student model, but instead distill the distribution of the original model to its counterpart vision-language distribution. Note that the loss is equivalent to the KL divergence loss between $o$ and the predicted distribution, since the additional entropy term $H(o)$ that appears in the KL divergence loss is a constant that does not depend on the MLP parameters. Eq. 1 shows that our approach is label-free and steers the MLP to be faithful to the original classifier $F$, since it is explicitly trained for that purpose. Note also that this loss function naturally encodes the classifier’s relationship across all classes.
We provide an illustration of the TexUnlock approach in Figure 1 and a PyTorch-like pseudocode in Section A of the Appendix. It is important to note that we only use the class name to formulate the textual prompt $l^{p}$, and no other supplementary information such as class descriptions, concepts or hierarchies (see Section O in the Appendix for more information).
After training, the projected visual features and the text encoder features lie in the same space. We can therefore query the visual features with any text by finding the alignment score between the embedded text and the projected visual features. In the case of image classification, the text queries remain the class prompts, and encoding them with the text encoder $T$ is equivalent to generating the weights of a linear classifier for the classification task formulated as argmax($\tilde{f}.U^{T}$), see Figure 1(b).
@@ 3.2 Zero-Shot Concept Bottleneck Models
Once the classifier’s distribution is matched to its corresponding vision–language counterpart distribution via TextUnlock, we proceed to formulate the proposed zero-shot CBMs. Note that at this stage, all model components (including the MLP) are frozen and no further training is performed. CBMs consist of two steps: (1) concept discovery, followed by (2) concept-to-class prediction. In step (1), the dense output features of a visual encoder are first mapped to textual concepts (e.g., words or short descriptions of objects) each with a score that represents the concept activation to the image. In step (2), a linear classifier $W^{con}$ is trained on top of these concept activations to predict the class.
Concept Discovery. We remind readers from Section 3 that $U\in\mathbb{R}^{K\times m}$ is the output of the text encoder $T$ for the class prompts, which in essence represent the classification weights of the newly formulated classifier. We assume that we are given a large set of predefined textual concepts, denoted as $\mathcal{Z}$, and with cardinality $|\mathcal{Z}|=Z$. Following other works (Rao2024DiscoverthenNameTC), and without loss of generality, we use the $Z=20K$ most common words in English (mostcommongoogle) as our concept set. These are general concepts that are sufficiently expressive and represent world knowledge and are not tailored towards any specific dataset. We ablate on other concept sets in Appendix Section I for the interested reader. To ensure that the concepts are meaningful, we apply a rigorous filtering procedure to the concept set. Specifically, we remove any terms that exactly match the target class name, as well as any constituent words that form the class name (for example, eliminating “tiger” and “shark” when the class name is “tiger shark”). In addition, we exclude terms corresponding to the parent and subparent classes (e.g., “fish” and “animal” for the class “tiger shark”), other species within the same category, and any synonyms of the target class name. Details of this procedure can be found in Section N of the Appendix. This systematic filtering guarantees that the resulting concept set is free of terms that are overly similar or directly derived from the target classes. With this filtering procedure, since no concept appears in the training data, our concept discovery set becomes entirely zero-shot. This will also demonstrate the MLP’s ability to generalize to the semantic space surrounding the class name.
Figure 2: Building zero-shot CBMs for any pretrained classifier. We first perform (a) concept discovery, followed by (b) building the concepts-to-class classifier in a zero-shot manner, which results in (c) our final CBM. Note that the concept bank only needs to be encoded once.
@@
Concept-to-class Prediction. The classifier $W^{con}$ takes in concept activations and outputs a distribution $S_{cn}$ over classes. We build $W^{con}$ in a zero-shot manner, where here zero-shot indicates that no training is required to map concept activations to classes. Recall that both $U$ and $C$ are outputs of the text encoder $T$, and they are already in the same space. Therefore, we can build the weights of the classifier $W^{con}$ with a text-to-text search between the concepts and the class name. Specifically, we calculate the cosine similarity between the concept embeddings $C$ and the classification matrix $U$ to obtain the new weights for $W^{con}$. That is, we perform $C\cdot U^{T}\in\mathbb{R}^{Z\times K}$. Therefore, the weights of $W^{con}$ represent how similar the class name is to each of the concepts. This process is shown in Figure 2(b). In total, the output distribution $S_{cn}$ of the CBM is defined as:
@@
From Eq. 2 we make an interesting observation. Our formulation involves scaling the linear feature-based classifier $U$ by the gram matrix of concepts ($C^{T}C\in\mathbb{R}^{m\times m}$). The gram matrix represents a feature correlation matrix measuring how different dimensions of the feature space relate to each other. Notably, if the gram matrix is the identity ($C^{T}C=I$), we get back our original feature-based classifier given by $\tilde{f}.U^{T}$. Therefore, to convert any classifier to a CBM, we plug in the gram matrix in-between, making it a convenient way to directly switch to an inherently interpretable model. Eq. 2 also shows that we do not change the linear classifier $U$, we only scale it by the gram matrix of concepts. This means our CBMs preserve the basic reasoning process of the original classifier. By this, we obtain zero-shot CBMs that discover concepts and build $W^{con}$ in a zero-shot manner for any classifier (Figure 2(c)). An additional unique property is the construction of CBMs at inference time, allowing any concept set to be selected to build a CBM on-the-fly. This makes our method highly flexible with respect to the chosen concept set.